OpenStudy (anonymous):

4.The total leverage of a business is given by the formula T= L+D I where T = total leverage, L = liabilities, D = debts, and I = income. A business can use the inequality T> L+D I to increase their buying power. Rewrite this formula to solve for D. Show all steps in your work.

OpenStudy (agent0smith):

T = L + DI is that the formula...?

OpenStudy (anonymous):

Yes

OpenStudy (agent0smith):

T = L + DI Rewrite this formula to solve for D. You want to get D by itself. Start by subtracting L from both sides of: T = L + DI

OpenStudy (anonymous):

My partner is doing that one so nvm. I need help on this one tho

OpenStudy (anonymous):

lever acts like a fulcrum to help lift heavy objects. The formula for the placement of the fulcrum is f1 · d1 = f2 · d2 where f1 = force applied at one end, d1 = distance of the first force from the fulcrum, f2 = force applied at the other end, and d2 = distance of the second force from the fulcrum. Rewrite this formula to solve for d1. Show all steps in your work.

OpenStudy (agent0smith):

\[\Large f_1 * d_1 = f_2 *d_2\] Rewrite this formula to solve for d1. There's only one step here to get d1 alone: divide both sides by f1

OpenStudy (anonymous):

So the answer is D1= (f2*D2/F1) ?

OpenStudy (agent0smith):

Correct :D

OpenStudy (anonymous):

Oh that was easy cx Thank you (:

OpenStudy (anonymous):

Can you help me with one more ?

OpenStudy (agent0smith):

sure

OpenStudy (anonymous):

2.Ohm’s Law states that the voltage across a wire is given by the formula V = IR where V = voltage, I = amperage, and R = resistance. Solve V = IR for R and use your formula to determine the resistance needed to produce 110 volts at 10 amps. Show all steps in your work.

OpenStudy (anonymous):

@sherimontoya @Euler271

OpenStudy (agent0smith):

V = IR Solve R How could you get R by itself? (what did we do with the last formula...)

OpenStudy (anonymous):

multiply by v ... ?

OpenStudy (agent0smith):

That won't work... V is already on the other side of the equals sign. We need to get R on it's own, ie get rid of the I V = I*R Think what we did in the last one... it'll be almost the same thing here.

OpenStudy (anonymous):

so we multiply I and r ?

OpenStudy (agent0smith):

No, they already are multiplied together (that's what IR means... I*R)

OpenStudy (agent0smith):

V = I*R is the same as V = IR you need to get R on it's own, by getting rid of the I in front of it.

OpenStudy (agent0smith):

Remember the last one... f1∗d1=f2∗d2 you had to divide both sides by f1 to get d1 by itself. You'll need to do basically the same thing with V = I*R

OpenStudy (agent0smith):

@lexis98 understand...?

OpenStudy (anonymous):

Kind of. But I do not understand what I am suppose to divide by

OpenStudy (agent0smith):

Well we have V = I*R... so I is being multiplied by R - how do you undo multiplication? division. so to get rid of the I, divide both sides by it.

OpenStudy (anonymous):

so v/I=r/I ?

OpenStudy (agent0smith):

Close. When you divide I*R by I, the I's cancel off leaving R The left side is correct.

OpenStudy (anonymous):

So the answer is V=R ? Idk..